Show that P1 (x) is orthogonal to [Pn (x)] ^2
on the interval (1,-1)
1
Expert's answer
2020-02-25T11:27:56-0500

We need to prove that "P_1(x)" is orthogonal to "(P_n(x))^2" on the interval "(-1,1)" "\Leftrightarrow" "\int\limits_{-1}^{1}P_1(x)(P_n(x))^2dx=0"

Proof:

"P_1(x)=x" is an odd function on "(-1,1)" .


One of the properties of Legendre polynomials is

"P_n(-x)=(-1)^nP_n(x)"

Using this property, we have

"(P_n(-x))^2=((-1)^nP_n(x))^2=(-1)^{2n}(P_n(x))^2=(P_n(x))^2"

So, "(P_n(x))^2" is an even function on "(-1,1)"


Therefore, "P_1(x)(P_n(x))^2" is an odd function on "(-1,1)" .


We can use the fact that

"f(x)" is odd "\Rightarrow \int\limits_{-a}^{a}f(x)dx=0"

for function "P_1(x)(P_n(x))^2" .


"P_1(x)(P_n(x))^2" is odd "\Rightarrow \int\limits_{-1}^{1} P_1(x)(P_n(x))^2 dx=0" .